# Enhancing Monotonic Modeling with Spatio-Temporal Adaptive Awareness in Diverse Marketing

### v1

@@ Abstract.
Monotonic Modeling; Online Food Ordering; Marketing Policy
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 744ccs: Information systems Location based services
@@ 1. Introduction
Users exhibit different response patterns to platform incentives across various marketing campaigns. Overall, the user Conversion Rate (CVR) expresses a monotonic increasing trend to the increment of red packets or discount amounts, while it monotonically decreases about the threshold for coupon. Similarly, CVR exhibits a monotonic decreasing relationship with delivery fees.
@@ 3. Preliminaries
(5) $-\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}r_{i,j}+\sum_{i=1}^{|\mathcal{C}|}(\mu_{i}(\sum^{|\mathcal{T}|}_{j=1}x_{i,j}-1))$
@@
(6) $\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}[r_{i,j}(\lambda t_{j}-\lambda\mathcal{B}-1)]$
@@ 4. Methods
In our feature extraction process, we segment the period into five intervals: breakfast [04:00–10:00), lunch [10:00–14:00), afternoon tea [14:00–17:00), dinner [17:00–20:00), and midnight snack [20:00–04:00), and we statistic numerical features to each period. We construct temporal features by incorporating information such as the Gregorian and lunar dates, holiday occurrences, and weekdays for each day. Simultaneously, we formulate spatial features from geographical data, encompassing the user’s current location, city ID, district ID, and AOI (Area of Interest) ID. Notably, in certain marketing campaigns different from user-level modeling, we employ regional modeling for AOIs. Therefore, depending on marketing campaigns, we tailor our features to use either users or AOIs (e.g., cross features, historical click sequences).
Spatio-temporal Correlation Activation. Within the Temporal Activation module, spatio-temporal features input to compute the weights of each feature embedding (i.e., User, Shop, Query) $e_{l_{i}}$ through a dot-product operation followed by a sigmoid function. This process aims to capture the dynamic variations inherent in spatio-temporal representations. The outcome is then multiplied by the input representations and integrated via a residual connection, designed to activate the spatio-temporal information encapsulated within the feature embeddings. The output of the Temporal Activation module, denoted as $h_{ta}$, can be formulated as follows:
@@
Employing the Temporal Target Attention mechanism to catch the interest preferences of different AOIs/Users towards Shops while simultaneously learning the variation across different times and spaces effectively. We achieve this by the linear projection of the Query, Key, and Value vectors with weights and bias (e.g., $[w_{Q},b_{Q}]$) generated by a fully connected neural network respectively. The module output denoted as $h_{tta}$, is articulated as follows:
@@
And we define the concave activation function $\hat{\rho}(x)$ and the saturated activation function $\tilde{\rho}(x)$ as follows:
@@
Simultaneously, to flexibly adapt to the spatiotemporally diverse property of marketing campaigns and learn the monotonic response functions, we introduce a strictly monotonic convex activation function, named Convex Linear Unit (CLU). Furthermore, we derive the corresponding concave activation function $\hat{\rho}(x)$ and the saturated activation function $\tilde{\rho}(x)$ through CLU. The activation functions construction is depicted in Figure 3, with CLU mathematical expressions as follows:
@@
The Universal Approximation Theorem(Cybenko, 1989; Hornik, 1991; Pinkus, 1999) posits that any continuous function over a closed interval can be approximated by neural networks with nonpolynomial activation functions. Concurrently, any multivariate continuous monotonic function on compact subsets of $\mathbb{R}^{k}$ can be approximated by monotonic constrained neural networks with at most $k$ hidden layers using sigmoid activation(Daniels and Velikova, 2010). Inspired by this, we propose the Four-Parameter Model (FPM) through an adaptive translation of the sigmoid function, combined with the CMNN framework, to flexibly approximate monotonic response functions in complex scenarios.
Figure 2(a) elucidates the functional form and parametric behavior of the FPM method, we learn and interpret the incentive sensitivity curve through four parameters, each bearing tangible significance in marketing. The parameter $\omega_{0}$ represents the natural conversion rate in the absence of any platform-driven incentive. The parameter $\omega_{1}$ denotes the sensitivity of the user at the motivational inflection point $\omega_{2}$, while $\omega_{3}$ corresponds to the maximal conversion rate attainable by the user.
@@
(a) Amount
(b) Threshold
Figure 4. Incentive sensitivity monotonic response curve.
@@ 5.1. Experimental Setup
Dataset Description. We conducted experiments on real-world datasets from two marketing campaigns on Ele.me, one of the largest OFOS platforms in China. The training datasets span one week, while the test dataset covers a single day. Notably, we randomly select 5% of online traffic flow as the detect traffic to randomly allocate incentive values as the ground truth for the incentive sensitivity function approximation. i) Exploding Red Packets Dataset: Send an average of 10 million exploding red packets to users online per day. It’s a classification dataset with approximately 1.05 million daily training samples and nearly 400 features, ultimately reaching 8.4 million samples. To evaluate the predictive accuracy of sensitivity functions, we trained models using unbiased samples collected from online detect traffic. ii) Delivery Fee Waiver Dataset: It is a regression dataset covering approximately 3.2 million AOIs, with about 4.3 million shops and 303 features. The daily samples are around 770 million, with the total exceeding 6 billion samples.
@@
Evaluation Metric. To comprehensively evaluate the performance of our model, we assess it from both accuracy and distribution perspectives. We first employ metrics such as the Area Under the Curve (AUC)(Fawcett, 2006), Mean Absolute Error (MAE)(Willmott and Matsuura, 2005) and Mean Squared Error (MSE)(Hyndman and Koehler, 2006) to evaluate the precision of the model. Subsequently, we use the Kullback-Leibler Divergence (KL Div)(Kullback and Leibler, 1951) and Correlation Coefficients (Corr Coeff)(Pearson, 1895) to assess the similarity between the model’s predicted incentive sensitivity functions and the detect traffic (i.e. ground truth).
@@
Baselines. To verify the effectiveness of our proposed methodology, we compare it against baselines with several state-of-the-art methods as follows: a) DNN: This is a naive model devoid of spatio-temporal and monotonic modeling capabilities. b) DNN-M: This method constrains the neural network weights to be exclusively non-negative or non-positive and combines sigmoid activation functions for monotonic modeling. c) FPM: This approach utilizes the mathematical four-parameter function to independently approximate the monotonic response function. d) CMNN: This method employs the CMNN framework in conjunction with different base convex activation functions, including CLU proposed in our study, while combining sigmoid as the final activation function.
@@
Hyperparamters. In our experiments, for Exploding Red Packets, the model training batch size is set to 128 with Adagrad as an optimizer and a learning rate of 0.001. In Delivery Fee Waiver, the batch size is configured to 1024, using an Adagrad optimizer with a learning rate of 0.012. Furthermore, all experiments are conducted on AOP [^4] A model training platform self-developed by the Alibaba Group using 40 parameter servers and 400 worker threads.
@@ 5.2. Experimental Results
Ablation Study To comprehensively understand the capability of our proposed CoMAN model in approximating incentive sensitivity functions, and recognizing that the S-t attention and adaptive activation modules constitute the core modeling components of CoMAN, we carried out exhaustive ablation studies to evaluate the effectiveness of these pivotal modules. Accordingly, we introduced three ablation variants: a) CoMAN w/o (S-t and AA) (CoMAN-B), this baseline version employs CMNN based on CLU combined with FPM as the final activation function. b) CoMAN w/o AA, this revision excludes the Adaptive Activation module from the CoMAN model. c) CoMAN w/o S-t, this variant omits the S-t Attention module from the CoMAN structure. These versions are evaluated against the complete CoMAN model, which integrates all the modules, to systematically assess the impact of each module on the overall model performance. Specifically, the comparative experiments of CMNN without the spatio-temporal activation and the spatio-temporal target attention modules.
Meanwhile, we also utilize the CMNN framework to demonstrate the effectiveness of the proposed CLU activation function in adaptively adjusting incentive responses in diverse marketing campaigns. To validate the monotonic response property of the CoMAN, we conduct offline tests in which all non-monotonic features are held constant, while the monotonic features (e.g., amount and threshold) increase. Then construct the CVR monotonic response curves, as illustrated in Figures 4(a) and 4(b) respectively. To evaluate our monotonic modeling task, we present the output of the monotonic layer from offline experiments in Table 1, while Table 2 and Table 3 showcase the performance of the model during online marketing campaigns.
@@
The results illustrated in the above tables indicate that the two modules within the CoMAN model’s monotonic layer align with our expectations, achieving significant improvements in both accuracy and prediction of real distributions. This demonstrates that CoMAN can adaptively adjust the convexity, concavity, and expression of the monotonic response functions in diverse marketing by enhancing the perception of spatio-temporal information. Moreover, as observed in Figure 5, we evaluate the model’s online predicted monotonic response curves both overall and across five periods, as well as for the top three cities with the most users. The fitting results for both the ablated and complete versions of CoMAN more closely approximate the ground truth distribution compared to the baseline, further verifying CoMAN’s effective capture of spatio-temporal characteristics in real-world marketing campaigns.

### v2

@@ Abstract.
Monotonic Modeling; Online Food Ordering; Marketing Policy; Budget Allocation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 168ccs: Information systems Location based services
@@ 1. Introduction
Users exhibit different monotone response patterns to platform incentives across various marketing campaigns(Yu et al., 2021; Zhao et al., 2019). It is reasonable that a higher incentive amount increases the probability of user redemption. Specifically, the user Conversion Rate (CVR) expresses a monotonic increasing trend to the increment of red packets or discount amounts, while it monotonically decreases about the threshold for coupon. Similarly, it exhibits a monotonic decreasing relationship with delivery fees. Without incorporating the monotone prior into the predictive model of user response, the model is susceptible to learning biased and over-fit patterns influenced by the noise introduced in the dataset.
@@ 3. Preliminaries
(5) $\min_{x_{i,j}}$ $-\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}r_{i,j}+\sum_{i=1}^{|\mathcal{C}|}(\mu_{i}(\sum^{|\mathcal{T}|}_{j=1}x_{i,j}-1))$
@@
(6) $\min_{x_{i,j}}$ $\sum_{i=1}^{|\mathcal{C}|}\sum_{j=1}^{|\mathcal{T}|}x_{i,j}[r_{i,j}(\lambda t_{j}-\lambda\mathcal{B}-1)]$
@@ 4. Methods
In our feature extraction process, we segment the period into five intervals: breakfast [04:00–10:00), lunch [10:00–14:00), afternoon tea [14:00–17:00), dinner [17:00–20:00), and midnight snack [20:00–04:00), and we statistic numerical features to each period. We construct temporal features by incorporating information such as the Gregorian and lunar dates, holiday occurrences, and weekdays for each day. Simultaneously, we formulate spatial features from geographical data, encompassing the user’s current location, city ID, district ID, and AOI (Area of Interest) ID. We utilize embedding matrices to semantically encode user attributes, shop attributes, queries, user behaviors, and spatio-temporal attributes respectively. Specifically, the latent embeddings $e_{l_{i}}$ for attribute features are derived from $[e_{l_{o}},e_{l_{u}},e_{l_{s}}]$, which correspond to cross features embedding, user embedding, and shop embedding, respectively. The query embedding is denoted as $e_{q}$, the behavior embedding as $e_{b}$, and the spatio-temporal embedding as $e_{st}$. Notably, in certain marketing campaigns different from user-level modeling, we employ regional modeling for AOIs. Therefore, depending on marketing campaigns, we tailor our features to use either users or AOIs (e.g., cross features, historical click sequences).
Spatio-temporal Correlation. We design two spatio-temporal correlation modules, namely Temporal Activation and Temporal Target Attention, to respectively explore the spatio-temporal characteristics and preferences inherent in static attributes and user behaviors. Within the Temporal Activation module, spatio-temporal features input to compute the weights of each feature embedding (i.e., User, Shop, Query) $e_{l_{i}}$ through a dot-product operation followed by a sigmoid function. This process aims to capture the dynamic variations inherent in spatio-temporal representations. The outcome is then multiplied by the input representations and integrated via a residual connection, designed to activate the spatio-temporal information encapsulated within the feature embeddings. The output of the Temporal Activation module, denoted as $h_{ta}$, can be formulated as follows:
@@
Employing the Temporal Target Attention mechanism to catch the interest preferences of different AOIs/Users towards Shops while simultaneously learning the variation across different times and spaces effectively. We accomplish this by applying linear projections to the Query, Key, and Value vectors using weights and biases (e.g., $[w_{Q},b_{Q}]$) generated by fully connected neural networks from $e_{st}$. The module output denoted as $h_{tta}$, is articulated as follows:
@@
And we define the concave activation function $\hat{\rho}(x)$ and the saturated function $\tilde{\rho}(x)$ as follows:
@@
Moreover, since the user response to incentives is not smooth across the entire treatment interval, it is impractical to model these curves using a linear representation with a fixed smoothing exponential function, as noted by (Zhao et al., 2019). Accordingly, to flexibly adapt to the spatiotemporally diverse property of marketing campaigns and learn the monotonic response functions, we introduce a strictly monotonic convex activation function, named Convex Linear Unit (CLU). The detailed proof of the convexity properties of CLU is provided in Appendix B.2. Furthermore, we derive the corresponding concave activation function $\hat{\rho}(x)$ and the saturated activation function $\tilde{\rho}(x)$ through CLU. The activation functions construction is depicted in Figure 3, with CLU mathematical expressions as follows:
@@
The Universal Approximation Theorem(Cybenko, 1989; Hornik, 1991; Pinkus, 1999) posits that any continuous function over a closed interval can be approximated by neural networks with nonpolynomial activation functions. Concurrently, any multivariate continuous monotonic function on compact subsets of $\mathbb{R}^{k}$ can be approximated by monotonic constrained neural networks with at most $k$ hidden layers using sigmoid activation(Daniels and Velikova, 2010). Inspired by this, we propose the Four-Parameter Model (FPM) through an adaptive translation of the sigmoid, combined with the CMNN framework, to flexibly approximate monotonic response functions in complex scenarios. Meanwhile, we derive Theorem 1, with a detailed proof available in the Appendix B.1.
Theorem 1.
Let $\breve{\rho}\in\mathcal{\breve{A}}$. Then any multivariate continuous monotone function f on a compact subset of $\mathbb{R}^{k}$ can be approximated with a monotone constrained neural network of at most $k$ layers using $\rho$ as the activation function.
Figure 2(a) elucidates the functional form and parametric behavior of the FPM method, we learn and interpret the incentive sensitivity curve through four parameters, each bearing tangible significance in marketing. The parameter $\omega_{0}$ represents the natural conversion rate in the absence of any incentive. The parameter $\omega_{1}$ denotes the sensitivity of the user at the motivational inflection point $\omega_{2}$, while $\omega_{3}$ corresponds to the maximal conversion rate attainable by users.
@@
Figure 4. Visualization of the average redemption amount of coupons in the top five cities with the most users across different periods.
@@ 5.1. Experimental Setup
Dataset Description. We conducted experiments on real-world datasets [^4] We gathered two real-world datasets from Ele.me’s online platform due to the absence of available public datasets suitable for evaluating our task. from two marketing campaigns on Ele.me, one of the largest OFOS platforms in China. The training datasets span one week, while the test dataset covers a single day. Notably, we randomly select 5% of online traffic flow as unbiased sample to randomly allocate incentive values as the ground truth for the incentive sensitivity function approximation. (i) Exploding Red Packets Dataset: Send an average of 10 million exploding red packets to users online per day. It’s a classification dataset with approximately 1.05 million daily training samples and nearly 400 features, ultimately reaching 8.4 million samples. (ii) Delivery Fee Waiver Dataset: It is a regression dataset covering approximately 3.2 million AOIs, with about 4.3 million shops and 303 features. The daily samples are around 770 million, with the total exceeding 6 billion samples. We segment application data by period and city and visualize users’ average redemption amount in exploding red packets, as depicted in Figure 4. The visualization reveals that users’ preferences for coupon amounts differ across various spatio-temporal settings, which is consistent with the analysis.
@@
Evaluation Metric. To comprehensively evaluate the performance of our model, we assess it from both accuracy and distribution perspectives. We first employ metrics such as the Area Under the Curve (AUC)(Fawcett, 2006), Mean Absolute Error (MAE)(Willmott and Matsuura, 2005) and Mean Squared Error (MSE)(Hyndman and Koehler, 2006) to evaluate the precision of the model. Subsequently, we use the Kullback-Leibler Divergence (KL Div)(Kullback and Leibler, 1951) and Correlation Coefficients (Corr Coeff)(Pearson, 1895) to assess the similarity between the model’s predicted incentive sensitivity functions and the random traffic (i.e. ground truth).
@@
Baselines. To verify the effectiveness of our proposed methodology, we compare it against baselines with several state-of-the-art methods as follows: (i) DNN: This is a naive model devoid of spatio-temporal and monotonic modeling capabilities. (ii) DNN-M: This method constrains the network weights to be exclusively non-negative or non-positive and combines sigmoid activation functions for monotonic modeling. (iii) SBBM: The Semi-black-box model(Zhao et al., 2019) predicts dynamic marketing responses by enhancing the logit demand curve through neural networks. (iv) FPM: This approach utilizes the mathematical four-parameter function to independently approximate the monotonic response function. (v) CMNN: This method employs the CMNN framework in conjunction with different base convex activation functions, including CLU proposed in our study, while combining sigmoid as the final activation function.
@@
Hyperparamters. In our experiments, for Exploding Red Packets, the model training batch size is set to 128 with Adagrad as an optimizer and a learning rate of 0.001. In Delivery Fee Waiver, the batch size is configured to 1024, using an Adagrad optimizer with a learning rate of 0.012. Furthermore, all experiments are conducted on AOP [^5] A model training platform self-developed by the Alibaba Group using 40 parameter servers and 400 worker threads. Further implementation details concerning the model parameters can be found in Appendix C.
(a) Exploding Red Packets
(b) Delivery Fee Waiver
Figure 6. Comparisons of uplift gain for users vary with different incentive sensitivity levels.
@@ 5.2. Experimental Results
Performance Comparison We report the overall performance of various competitive baselines and our proposed model across two datasets in Table 1. Moreover, to evaluate the real-world effectiveness of our approach in practical applications, we present the performance of the model in online marketing campaigns as shown in Table 2 and Table 3. The results yield the following observations: (i) The performance of CoMAN is notably superior to all competitive baselines. Compared to previous marketing response approaches, CoMAN achieves an average improvement of 5.33% in user redemption response accuracy and an average enhancement of 5.89% in distribution similarity. These remarkable performance gains strongly indicate that CoMAN has the potential to enhance marketing efficiency in budget allocation significantly. (ii) Methods like FPM and SBBM, which employ nonlinear mathematical functions in conjunction with neural networks to adaptively enhance the logit response curve, are more effective than simply applying positive or negative constraint learning to neural network weights. These methods are capable of more accurately characterizing users’ sensitivity to incentives. Specifically, FPM is more effective than SBBM in representing this sensitivity, thereby highlighting the effectiveness of our approach in integrating FPM for learning monotonic response. (iii) Compared to traditional marketing approaches like DNN-M and SBBM, CoMAN performs better in perceiving and incorporating spatio-temporal information. Furthermore, our model surpasses previous monotonic methods such as CMNN and FPM in effectively capturing spatio-temporal representations which enhances monotonic modeling. These findings underscore the significant advancements of CoMAN, primarily attributed to the activation modeling of spatio-temporal information and the robust capture of spatio-temporal monotonic patterns. (iv) Our designed CLU function, when integrated with CMNN, outperforms the other two activation functions, ReLU and ELU. This outcome substantiates the validity of approximating user monotonic response curves with a non-fixed exponential function, thus better capturing and expressing user sensitivity to incentives. (v) Our approach exhibits superior performance across datasets with different modeling granularities (i.e., user-level and regional-level). This indicates that our method is capable of more precisely describing monotonic response under various incentives, whether the focus is on user-level perception or regional-level modeling.
(a) Amount
(b) Threshold
Figure 7. Monotonic response curves of user sensitivity to incentives with different amounts and thresholds
To visualize the effectiveness of our model in budget allocation compared to previous marketing methods, we follow the idea proposed by (Yu et al., 2021; Liu et al., 2019) and use the first derivative of the response curve at the inflection point as the ”gradient” representing a user’s sensitivity to incentives. Under the same incentive treatment, user groups with high ”gradient” values yield greater commercial returns compared to those with low ”gradient”. Therefore, the efficiency of the marketing model can be attributed to the accurate estimation of response scores for users with varying incentive sensitivities. Subsequently, we rank the users in descending order based on their ”gradient” and divide them into five groups representing different levels of sensitivity to incentives. As illustrated in Figure 6, CoMAN predicts a higher uplift gain for the first group of highly sensitive users compared to SBBM and FPM, while predicting a correspondingly lower uplift gain for the least sensitive users. This demonstrates our model’s higher efficiency in budget allocation.
@@
Ablation Study To comprehensively understand the capability of our proposed CoMAN model in approximating incentive sensitivity functions, and recognizing that the S-t attention and adaptive activation modules constitute the core modeling components of CoMAN, we carried out exhaustive ablation studies to evaluate the effectiveness of these pivotal modules. Accordingly, we introduced three ablation variants: (i) CoMAN w/o (S-t and AA) (CoMAN-B), this baseline version employs CMNN based on CLU combined with FPM as the final activation function. (ii) CoMAN w/o AA, this revision excludes the Adaptive Activation module from the CoMAN model. (iii) CoMAN w/o S-t, this variant omits the S-t Attention module from the CoMAN structure. These versions are evaluated against the complete CoMAN model, which integrates all the modules, to systematically assess the impact of each module on the overall model performance. Additionally, the comparative experiments of CMNN without the spatio-temporal activation and the spatio-temporal target attention modules.
To validate the monotonic response property of the CoMAN, we conduct offline tests in which all non-monotonic features are held constant, while the monotonic features (e.g., amount and threshold) increase. Then construct the CVR monotonic response curves, as illustrated in Figures 7(a) and 7(b) respectively. From the results shown in the above tables, we have the following observations: (i) As our analysis suggests, models that leverage spatio-temporal information in the monotonic layer to enhance the perception of the monotonic response outperform models that consider only basic attributes and behavioral representations (e.g., FPM, CMNN) or incorporate spatio-temporal capture solely in the embedding (e.g., CoMAN-B). (ii) The performance of single-module models in the monotonic layer is significantly lower than that of models employing a two-module perception framework. This finding is consistent with our design rationale, indicating that two-module perception of spatio-temporal information more effectively captures the additional semantics inherent in user monotonic responses. (iii) As anticipated, employing the Adaptive Activation module facilitates more adaptive learning of the concavity and convexity proportions of the response function by harnessing spatio-temporal information. Concurrently, utilizing the S-t Attention module enables a more precise spatio-temporal representation of key parameters in the sensitive response curves.
@@
Moreover, as observed in Figure 5, we evaluate the model’s online predicted monotonic response curves both overall and across five periods, as well as for the top three cities with the most users. The fitting results for both the ablated and complete versions of CoMAN more closely approximate the ground truth distribution compared to the baseline, further verifying CoMAN’s effective capture of spatio-temporal characteristics in real-world marketing campaigns.